OpenStudy (anonymous):

Simplify the equation and write it in standard form. 168=(x+1)(x-4)(x)

OpenStudy (anonymous):

@Baby_Bear69

OpenStudy (vheah):

1) Multiply: Distribute the (x) to (x+1)(x-4) and you will get \[(x ^{2}+x)(x ^{2}-4x)\] 2) Foil that equation to get one whole answer, instead of two things being multiplied to each other.

OpenStudy (anonymous):

You distribute (x) to (x+1) or to (x-4), but no to both. if you do so you would end with a quartic equation instead of a cubic equation.

OpenStudy (anonymous):

You can also foil (x+1)(x-4) = x^2-3x-4 and then distribute (x) to that to get x^3-3x^2-4x. then subtract 168 from both sides and you end with x^3-3x^2-4x-168

Directrix (directrix):

x^3 -3x^2 -4x - 168 = 0

OpenStudy (anonymous):

@directrix is that the final answer?

Directrix (directrix):

Simplify the equation and write it in standard form. The polynomial in standard form is: x^3-3x^2-4x-168 But, the instructions ask for the polynomial equation in standard form. I assume that to be this: f(x) = x^3-3x^2-4x-168 @j_slate233

OpenStudy (anonymous):

Thank you!!! @Directrix

Directrix (directrix):

You are welcome.